A race car is about to pass a curve road with a diameter of 450m at a constant speed of 22.5m/s. What is the centripetal acceleration of the race car?





1
Expert's answer
2022-02-28T10:40:34-0500
"a=\frac{v^2}{r}=\frac{22.5^2}{450}=1.125\:\rm m/s^2"